# On the Theory of Risk-Aware Agents: Bridging Actor-Critic and Economics

### v2

@@ 1 Introduction
Figure 1: We assess the proposed method (DAC) on a range of DeepMind Control and MetaWorld tasks listed in Table 2. We compare DAC to a variety of model-free algorithms, in both low and high replay regimes. In high replay ratio, all algorithms are scaled by resetting. Our findings indicate that DAC achieves significant improvements in performance and sample efficiency compared to the Soft Actor-Critic. The $Y$-axis denotes the interquartile mean and the $95$% confidence interval. $21$ tasks, $1$mln environment steps, and 10 seeds per task.
@@
The attitudes towards risk of algorithmic agents have been researched in multiple contexts. For instance, a risk-seeking approach of optimism in the face of uncertainty has been identified as an effective exploration strategy, minimizing regret during the learning process (Wang et al., 2020b; Neu & Pike-Burke, 2020). Conversely, risk-averse, pessimistic learning strategies have proven beneficial in counteracting value overestimation caused by temporal difference errors (Hasselt, 2010; Fujimoto et al., 2018). However, there is a disconnect between these risk-aware strategies and the foundational theories of RL, particularly in how they relate to the goal of value maximization. As a result, despite the empirical success of risk-aware agents, the specific class of policies they implement remains ambiguous.
@@
In this paper, we study the reinforcement learning objective aligned with the principles of decision theory: the Von Neumann–Morgenstern theorem (Von Neumann & Morgenstern, 1947); and the expected utility hypothesis (Von Neumann & Morgenstern, 1947; Kahneman & Tversky, 1979). We show that, in contrast to pure value maximization, the proposed theory allows for derivation of both risk-neutral (e.g., DDPG (Silver et al., 2014)) and risk-aware approaches (e.g., TD3 (Fujimoto et al., 2018)). We demonstrate that the pessimistic updates in state of the art algorithms such as SAC (Haarnoja et al., 2018), REDQ (Chen et al., 2020), or TOP (Moskovitz et al., 2021) can be derived from expected utility maximization under an exponential utility function. Furthermore, we introduce Dual Actor-Critic (DAC), a risk-aware algorithm with two actors: optimistic and pessimistic. In DAC, each actor is trained independently using gradient backpropagation of a distinct objective: the optimistic actor aims to maximize an upper value bound for exploration, while the pessimistic actor focuses on a lower bound for temporal-difference (TD) targets (Fujimoto et al., 2018; Haarnoja et al., 2018). DAC also features an automatic adjustment mechanism to balance the divergence between these actors, allowing for adaptability to various tasks without hyperparameter tuning. We evaluate DAC on a diverse set of locomotion and manipulation tasks and find that, despite its simplicity, DAC achieves state of the art sample efficiency and performance. Below, we outline our contributions:
@@
$\mathbf{1}$. We introduce a generalized utility-based actor-critic objective, capable of formalizing both risk-neutral and risk-aware actor-critic algorithms. Based on the proposed theory, we demonstrate that the policies enacted by risk-aware algorithms align with the certainty equivalent of value, and as such are optimal in the decision theoretic context. We show that a policy optimizing Clipped Double Q-learning is approximately utility-optimal under an exponential utility.
@@
$\mathbf{3}$. We show that DAC outperforms benchmarks in terms of both sample efficiency and final performance, in both low and high replay regimes. Notably, DAC solves the dog domain, reaching the performance of significantly more complex model-based methods. To facilitate further research, we perform ablations on various design and hyperparameter choices (over $5000$ training runs). We release the implementation of DAC under the following URL.
@@ 2.1 Reinforcement Learning
We consider an infinite-horizon Markov Decision Process (MDP) (Puterman, 2014) which is described with a tuple $(S,A,r,p,\gamma)$, where states $S$ and actions $A$ are continuous, $r_{s,a}$ is the transition reward, $p(s)$ is the initial state distribution and $\gamma\in(0,1]$ is a discount factor. A policy $\pi(a|s)$ is a state-conditioned action distribution with entropy denoted as $\mathcal{H}(s)$. Soft Value is the sum of expected discounted return and state entropies from following the policy at a given state $V^{\pi}(s)=\mathrm{E}_{a\sim\pi}\left[r_{s,a}+\alpha\mathcal{H}(s)+\gamma V^{\pi}(s^{\prime})\right]$, with $\alpha$ denoting the entropy temperature parameter. Q-value is the expected discounted return from performing an action and following the policy thereafter $Q^{\pi}(s,a)=\left[r_{s,a}+\gamma V^{\pi}(s^{\prime})\right]$. A policy is said to be optimal if it maximizes the expected value of the possible starting states $s_{0}$, such that $\dot{\pi}=\arg\max_{\pi\in\Pi}\mathrm{E}_{s_{0}\sim p}V^{\pi}(s_{0})$, with $\dot{\pi}$ denoting the optimal policy and $\Pi$ denoting the considered set of policies (eg. Gaussian). In off-policy actor-critic, there is continuous gradient-based learning of both Q-values (critic) and the policy (actor). The critic parameters $\theta$ are updated by minimizing the variants of SARSA temporal-difference on transitions $T=(s,a,r,s^{\prime})$, with $T$ being sampled from past experiences (Fujimoto et al., 2018; Haarnoja et al., 2018):
[…]
Here, $Q_{\theta}(s,a)$ is the critic output, $V^{lb}_{\theta}$ is the bootstrap value derived from a target network, and $\mathcal{D}$ is the experience buffer (Mnih et al., 2015). Under certain conditions $Q_{\theta}(s,a)$ is guaranteed to perfectly approximate the Q-value under the policy from which the bootstrap is sampled (Van Seijen et al., 2009; Sutton & Barto, 2018). The policy parameters $\phi$ are updated to seek locally optimal values approximated by the critic (Ciosek & Whiteson, 2020):
$\begin{split}\dot{\pi}_{\phi}&=\underset{\pi\in\Pi}{\arg\max}\underset{s\sim\mathcal{D}}{\mathrm{E}}~{}V^{lb}_{\theta}(s)\end{split}$ (2)
In this equation, $\pi_{\phi}$ represents the actor, and $V^{lb}_{\theta}$ is the estimated lower-bound value, calculated using the critic:
$\begin{split}V^{lb}_{\theta}(s)\approx Q^{lb}_{\theta}(s,a)-\alpha\log\pi_{\phi}(a|s),\quad a\sim\pi_{\phi}(s)\end{split}$ (3)
@@
Where $Q^{lb}_{\theta}$ represents the lower bound Q-value approximation, $\log\pi_{\phi}$ is the state-action entropy. Employing $V^{lb}_{\theta}$, a technique called Clipped Double Q-Learning (CDQL), has proven effective in mitigating value overestimation in Temporal Difference (TD) learning (Van Hasselt et al., 2016; Fujimoto et al., 2018). The lower bound is derived from a critic ensemble, often with two networks (Fujimoto et al., 2018; Haarnoja et al., 2018; Ciosek et al., 2019):
@@
$\begin{split}Q^{lb}_{\theta}(s,a)&=\min\bigl{(}Q_{\theta}^{1}(s,a),Q_{\theta}^{2}(s,a)\bigr{)}\\ &=\underbrace{\frac{1}{2}\bigl{(}Q_{\theta}^{1}(.)+Q_{\theta}^{2}(.)\bigr{)}}_{\text{Mean}}-\underbrace{\frac{1}{2}~{}|Q_{\theta}^{1}(.)-Q_{\theta}^{2}(.)|}_{\text{Standard Deviation}}\end{split}$ (4)
The observation led to a generalization of the lower bound (Moskovitz et al., 2021; Cetin & Celiktutan, 2023):
$\begin{split}V^{\beta}_{\theta}(s)&\triangleq V^{\mu}_{\theta}(s)+\beta V^{\sigma}_{\theta}(s)\\ Q^{\beta}_{\theta}(s,a)&\triangleq Q^{\mu}_{\theta}(s,a)+\beta Q^{\sigma}_{\theta}(s,a)\end{split}$ (5)
In this formulation, $\beta$ represents the degree of pessimism, and $(\mu,\sigma)$ denote the mean and standard deviation of the value model ensemble and $V^{\beta}_{\theta}(s)=\mathrm{E}_{a\sim\pi}Q^{\beta}_{\theta}(s,a)-\log\pi_{\phi}(a|s)$. In this formulation, the risk level of the lower bound is modulated by $\beta$. In case of $\beta=-1$, the pessimistic value $Q^{\beta}_{\theta}$ equals the CDQL lower bound $Q^{lb}_{\theta}$. A non-zero $\beta$ indicates a departure from risk-neutrality: positive $\beta$ yields a pessimistic lower bound, while negative $\beta$ results in an optimistic upper bound. Optimizing actor-critic modules towards a pessimistic lower bound discourages state actions with high critic ensemble disagreement, a phenomenon known as pessimistic underexploration (Ciosek et al., 2019; Cetin & Celiktutan, 2023). Given fixed policy parameters, it follows that optimization towards the risk-aware bootstrap leads value approximator $V_{\theta}$ to an output:
$\begin{split}&V_{\theta}(s)=\underset{a_{i}\sim\pi}{\mathrm{E}}~{}\sum_{i=0}^{\infty}\gamma^{i}\bigl{(}r_{s_{i},a_{i}}+\alpha\mathcal{H}(s_{i})+\beta V^{\sigma}_{\theta}(s_{i})\bigr{)}\end{split}$ (6)
The subscript $i$ denotes the outcomes in states following state $s$. Thus, for given $\phi$ the difference between the risk-aware value (resulting from bootstrapping with $V^{\beta}_{\theta}$), and risk-neutral values (resulting from bootstrapping with $V^{\mu}_{\theta}$ and denoted as $\bar{V}_{\theta}$) is proportional to the expected values of the discounted sum of critic disagreements, as shown below:
$\begin{split}V_{\theta}(s)-\bar{V}_{\theta}(s)=\underset{a_{i}\sim\pi}{\mathrm{E}}~{}\beta\sum_{i=0}^{\infty}\gamma^{i}~{}V^{\sigma}_{\theta}(s_{i})\end{split}$ (7)
As such, if the difference is nonzero, the pessimistic policy diverges from the risk-neutral policy and yields regret. Despite this, numerous effective off-policy algorithms adopt a risk-aware objective, leaning towards either pessimism or optimism (Fujimoto et al., 2018; Haarnoja et al., 2018; Ciosek et al., 2019; Cetin & Celiktutan, 2023).
2.2 Expected Utility Theorem
The Von Neumann–Morgenstern Theorem (Von Neumann & Morgenstern, 1947) posits that an agent whose preferences adhere to four axioms (completeness, transitivity, continuity, and independence), has a utility function $\mathcal{U}(x)$ that function enables the comparison of the preferences. The expected utility hypothesis (Von Neumann & Morgenstern, 1947; Kahneman & Tversky, 1979) states that agents choose between risky options by comparing their expected utility, such that given $x^{i}\sim X^{i}$, it follows that $X^{1}\succeq X^{2}\iff\mathrm{E}~{}U(x^{1})\geq\mathrm{E}~{}U(x^{2})$. This implies optimization of utility, rather than its input (Stiglitz, 1997).
[…]

### v3

@@ 1 Introduction
Figure 1: We test the proposed approach (DAC) against various risk-aware and risk-neutral actor-critic baselines in $30$ tasks listed in Table 2. Due to a controlled experimental setup described in Section 5, the performance differences between algorithms stem solely from their risk-management. $Y$-axis reports IQM with $95$% CI calculated using 10 seeds, with $1.0$ representing the maximal score.
@@
The attitudes towards risk of algorithmic agents have been researched in multiple contexts. For instance, a risk-seeking approach of optimism in the face of uncertainty has been identified as an effective exploration strategy, minimizing regret during the learning process (Wang et al., 2020b; Neu & Pike-Burke, 2020). Conversely, risk-averse, pessimistic Q-learning strategies have proven beneficial in counteracting value overestimation caused by temporal difference errors Hasselt (2010); Fujimoto et al. (2018). However, there is a disconnect between these risk-aware strategies and the foundational theories of RL, particularly in how they relate to the goal of value maximization. As a result, despite the empirical success of risk-aware agents like TD3 (Fujimoto et al., 2018) or SAC (Haarnoja et al., 2018a), the specific class of policies they implement remains ambiguous.
@@
In this paper, we study the reinforcement learning objective aligned with the principles of decision theory. We show that, in contrast to pure value maximization, the decision theoretic perspective allows for derivation of both risk-neutral (e.g., DDPG (Silver et al., 2014)) and risk-aware approaches (e.g., TD3 (Fujimoto et al., 2018)). As such, we demonstrate that the pessimistic updates in state-of-the-art algorithms such as SAC (Haarnoja et al., 2018a), REDQ (Chen et al., 2020), or TOP Moskovitz et al. (2021) can be derived from expected utility maximization under an exponential utility function. Furthermore, we introduce Dual Actor-Critic (DAC), a risk-aware algorithm with two actors: optimistic and pessimistic. In DAC, each actor is trained independently using gradient backpropagation of a distinct objective: the optimistic actor aims to maximize an upper value bound for exploration, while the pessimistic actor focuses on a lower bound for temporal-difference (TD) targets (Fujimoto et al., 2018; Haarnoja et al., 2018a). DAC also features an automatic adjustment mechanism to balance the divergence between these actors, allowing for adaptability to various tasks without hyperparameter tuning. We evaluate DAC on a diverse set of locomotion and manipulation tasks and find that, despite its simplicity, DAC achieves significant improvements in sample efficiency and performance. Below, we outline our contributions:
@@
$\mathbf{1}$. We consider a generalized utility-based actor-critic objective, capable of formalizing both risk-neutral and risk-aware actor-critic algorithms. We demonstrate that the policies enacted by Clipped Double Q-Learning and generalizations thereof approximately align with the certainty equivalent of value under an exponential utility, and as such are optimal in the decision theoretic context.
@@
$\mathbf{3}$. We show that DAC outperforms benchmark algorithms in terms of both sample efficiency and final performance. Notably, DAC solves the dog domain, reaching the performance of significantly more complex model-based methods. To facilitate further research, we perform ablations on various design and hyperparameter choices (over $5000$ training runs). We release the implementation of DAC under the following URL
@@ 2 Background
We consider an infinite-horizon Markov Decision Process (MDP) (Puterman, 2014) which is described with a tuple $(S,A,r,p,\gamma)$, where states $S$ and actions $A$ are continuous, $r_{s,a}$ is the transition reward, $p$ is a deterministic transition mapping, with $p_{0}$ being the starting state distribution, and $\gamma\in(0,1]$ is a discount factor. A policy $\pi(a|s)$ is a state-conditioned action distribution with entropy denoted as $\mathcal{H}(\pi(s))$. Soft Value (Haarnoja et al., 2018a) is the sum of expected discounted return and state entropies from following the policy at a given state $V^{\pi}(s)=\mathrm{E}_{a\sim\pi}\left[r_{s,a}+\alpha\mathcal{H}(\pi(s))+\gamma V^{\pi}(s^{\prime})\right]$, with $\alpha$ denoting the entropy temperature parameter. Q-value is the expected discounted return from performing an action and following the policy thereafter $Q^{\pi}(s,a)=\left[r_{s,a}+\gamma V^{\pi}(s^{\prime})\right]$. A policy is said to be optimal if it maximizes the expected value of the possible starting states $s_{0}$, such that $\pi^{*}=\arg\max_{\pi\in\Pi}\mathrm{E}_{s_{0}\sim p_{0}}~{}V^{\pi}(s_{0})$, with $\pi^{*}$ denoting the optimal policy and $\Pi$ denoting the considered set of policies (eg. Gaussian). Off-policy actor-critic algorithms perform gradient-based learning of both Q-values (ie. critic or value model) and the policy (ie. actor). The critic parameters $\theta$ are updated by minimizing SARSA temporal-difference loss $\mathcal{L}_{\theta}$ on transitions $T=(s,a,r,s^{\prime})$ which are sampled from past experiences (Fujimoto et al., 2018; Haarnoja et al., 2018a) according to $\mathcal{L}_{\theta}=\mathrm{E}_{T\sim\mathcal{D}}~{}(Q_{\theta}(s,a)-r_{s,a}-\gamma V^{\pi_{\phi}}(s^{\prime}))^{2}$. Here, $Q_{\theta}(s,a)$ is the critic output, $V^{\pi}(s^{\prime})$ is the bootstrap value derived from a target network, and $\mathcal{D}$ is the experience buffer (Mnih et al., 2015). The policy parameters $\phi$ are updated to maximize values approximated by the critic (Ciosek & Whiteson, 2020): $\mathcal{L}_{\phi}=\mathrm{E}_{s\sim\mathcal{D}}~{}V^{\pi_{\phi}}(s)$. In this equation, $\pi_{\phi}$ represents the actor, and $V^{\pi_{\phi}}(s)$ is the state value under $\pi_{\phi}$.
Risk-Awareness in Actor-Critic
@@
In standard actor-critic the critic represents the expected future discounted return and the actor is optimized with respect to such critic (Lillicrap et al., 2015). In contrast, popular algorithms like SAC or TD3 learn a critic that represents the lower bound of Q-values (Fujimoto et al., 2018; Haarnoja et al., 2018b). Employing the lower-bound has proven effective in mitigating value overestimation in Temporal Difference (TD) learning (Van Hasselt et al., 2016; Fujimoto et al., 2018). A popular approach is Clipped Double Q-Learning (CDQL), where the lower-bound is calculated by taking the minimum value of an ensemble of critics, most often two (Fujimoto et al., 2018; Haarnoja et al., 2018b; Ciosek et al., 2019; Hansen et al., 2022): $V^{\pi_{\phi}}(s)\approx\min(Q_{\theta}^{1}(s,a),Q_{\theta}^{2}(s,a))-\log\pi_{\phi}(a|s)$, with $a\sim\pi_{\phi}(s)$. Here, $Q_{\theta}^{1}(s,a)$ and $Q_{\theta}^{2}(s,a)$ denote the first and the second critic in the ensemble, and $\log\pi_{\phi}$ is the state-action entropy. Furthermore, it was shown that using the minimum is equivalent to ensemble statistics (Ciosek et al., 2019):
@@
$\begin{split}\min\bigl{(}Q_{\theta}^{1}(s,a),Q_{\theta}^{2}(s,a)\bigr{)}=\underbrace{\frac{1}{2}\bigl{(}Q_{\theta}^{1}(s,a)+Q_{\theta}^{2}(s,a)\bigr{)}}_{\text{Mean}}-\underbrace{\frac{1}{2}~{}|Q_{\theta}^{1}(s,a)-Q_{\theta}^{2}(s,a)|}_{\text{Standard Deviation}}.\end{split}$ (1)
This led to generalizations of the lower-bound as to include varying levels of pessimism (Ciosek et al., 2019; Moskovitz et al., 2021):
$\begin{split}Q^{\beta}(s,a)=Q^{\mu}(s,a)+\beta Q^{\sigma}(s,a),\quad\text{and}\quad V^{\beta}(s)=V^{\mu}(s)+\beta V^{\sigma}(s)\end{split}$ (2)
Above, $Q^{\beta}(s,a)$ and $V^{\beta}(s)$ are the pessimistic Q-value and values respectively, $\beta$ represents the degree of pessimism, and $(\mu,\sigma)$ denote the mean and standard deviation of the critic model ensemble. Furthermore, the pessimistic values and Q-values are related by $V^{\beta}(s)=\mathrm{E}_{a\sim\pi}(Q^{\beta}(s,a)-\log\pi_{\theta}(a|s))$. In this setup, the risk level of the lower bound is modulated by $\beta$. As shown in Equation 1, in case of $\beta=-1$ it is true that $Q^{\beta}=\min(Q_{\theta}^{1}(s,a),Q_{\theta}^{2}(s,a))$. In setup of pessimistic learning the value targets are affected by the uncertainty measured via the value model ensemble disagreement. A non-zero $\beta$ indicates a departure from risk-neutrality: positive $\beta$ yields a pessimistic lower bound, while negative $\beta$ results in an optimistic upper bound. Thus, for given $\phi$ the difference between the risk-aware value (resulting from bootstrapping with $Q^{\beta}$), and risk-neutral values (resulting from bootstrapping with $Q^{\mu}$) is proportional to the expected values of the discounted sum of critic disagreements (Fujimoto et al., 2018). As such, it is clear that iterating the pessimistic objective does not result in the optimal policy as long as the critic disagreement is not equal to zero for all state-actions (Kumar et al., 2020). Despite this, numerous effective off-policy algorithms adopt a risk-aware objective, leaning towards either pessimism or optimism (Fujimoto et al., 2018; Ciosek et al., 2019; Cetin & Celiktutan, 2023). This understanding of risk-awareness is slightly different to the risk-awareness considered in the distributional RL (Bellemare et al., 2017; Dabney et al., 2018b, a) or the CVaR series (La & Ghavamzadeh, 2013; Hiraoka et al., 2019; Lim & Malik, 2022). There, the distribution of returns is explicitly modelled via a distribution and the risk-awareness is calculated with respect to the aleatoric uncertainty (Dabney et al., 2018a; Lim & Malik, 2022). In contrast, CDQL and associated techniques use regular critics and consider the total estimation uncertainty associated with using point-estimate to model Q-values (Fujimoto et al., 2018; Hiraoka et al., 2021).
Expected Utility Theorem